A 2.3 kg otter starts from rest at the top of a
muddy incline 90.5 cm long and slides down
to the bottom in 0.50 s.
What net external force acts on the otter
along the incline?
Answer in units of N
1
Expert's answer
2011-10-28T08:09:31-0400
Let's use the second Newton's law:
F = ma.
All we need is to calculate acceleration a. Why can do it in a such way:
Va = 0.905/0.50 = 1,81 m/s - average speed of an ottet.
We know that V(0) = 0. Also we& know that V(0.5/2) = V(0.25) = Va.
So, a = (V(0.25)-V(0))/0.25 = (Va-0)*4 = 4*Va = 4*1.81 = 7,24 m/s^2.
At last we obtain F = ma = 2.3*7.24 = 16,652 N.